OpenStudy (anonymous):

A die is rolled 12 times. Find the chance that the face with six spots appears once among the first 6 rolls, and once among the next 6 rolls.

OpenStudy (kropot72):

The chance of a particular number not coming up in a roll of the die is 5/6. The probability of 6 not appearing in 5 out of 6 rolls is \[(\frac{5}{6})^{5}\]

OpenStudy (kropot72):

The events 'the face with six spots appears once among the first 6 rolls' and 'the face with six spots appears once among the next 6 rolls' are independent. Therefore the probability of both occurring is the product of the individual probabilities: \[Required\ probability=(\frac{5}{6})^{5}\times (\frac{5}{6})^{5}=(\frac{5}{6})^{10}\]

OpenStudy (anonymous):

I understand the approach, but i have 2 small questions: 1. isn't supposed to be (5/6)^6 ? 2. isn't supposed to be at the end P(one-6) = 1-P(no-6)? Thanks!